How many different sets of 4 students can be chosen out of 17 qualified students to represent a school in a mathematics contest?
Answer by Alan3354(69443)

How many different sets of 4 students can be chosen out of 17 qualified students to represent a school in a mathematics contest?
===============
17*16*15*14 = 57120
-------
If all positions are equal, then ABC & D is the same as BAD & C
-----
--> divide by 4*3*2*1 = 24
---
57120/24 = 2380 sets